Consider an electric circuit with an inductance of 0.05 henry, a resistance of 20 ohms, a condenser of capacitance of 100 micro farads and an E = 100 volts. Find I and Q given the initial conditions Q = 0, I = 0 at t = 0


1
Expert's answer
2022-03-22T04:21:07-0400


Due to Kirchhoff's law:

"U_L+U_R+U_C=E" (1)

where

"U_L=L\frac{di}{dt}" ; "U_R=iR" ; "U_C=\frac1C\int idt" .

Let's differentiate (1):

"L\frac{d^2i}{dt^2}+R\frac{di}{dt}+\frac 1C i=0"

"\frac{d^2i}{dt^2}+\frac RL\frac{di}{dt}+\frac {1}{LC} i=0"

Characteristic equation:

"\lambda^2+\frac RL\lambda+\frac {1}{LC}=0"

"\frac RL=\frac {20}{0.05}=400"; "\frac {1}{LC}=\frac {1}{0.05\cdot 100\cdot 10^{-6}}=2\cdot10^5".

"\lambda^2+400\lambda+2\cdot 10^5=0"

"\lambda=-200\pm400j" , where "j=\sqrt{-1}" .

"i(t)=e^{-200t}(C_1 \cos {400t}+C_2 \sin {400t})".

"i|_{t=0}=C_1=0;"

"L\frac{di}{dt}|_{t=0}+Ri|_{t=0}+U_C|_{t=0}=E" (2)

"U_C|_{t=0}=\frac 1C Q|_{t=0}=0" ; (3)

"Ri|_{t=0}=0"; (4)

"\frac{di}{dt}=C_2e^{-200t}(-200\cdot\sin{400t}+400\cdot\cos{400t})";

"\frac{di}{dt}|_{t=0}=400\cdot C_2". (5)

Substitute (3), (4) and (5) into (2):

"L\cdot400\cdot C_2=E";

"C_2=\frac{E}{400\cdot L}=\frac{100}{400\cdot 0.05}=5"

"i(t)=5\cdot e^{-200t}\cdot \sin{400t}."

"Q(t)=\int5\cdot e^{-200t}\cdot \sin{400t} dt="

"-\frac{1}{200}e^{-200t}(\sin{400t}+2\cos{400t})+C_3";

"Q|_{t=0}=-\frac{1}{200}\cdot2+C_3=0";

"C_3=\frac{1}{100}"

"Q(t)=\frac{1}{200}[2-e^{-200t}(\sin{400t}+2\cos{400t})]" .

Answer:

"i(t)=5\cdot e^{-200t}\cdot \sin{400t}" ;

"Q(t)=\frac{1}{200}[2-e^{-200t}(\sin{400t}+2\cos{400t})]" .